Determine the velocity of sound in mercury, which has bulk modulus of 2.8×10^10 and the a density of 1.36×10⁴ kg/m³.


1
Expert's answer
2021-05-07T08:13:13-0400

V( velocity )= "\sqrt { \frac{B}{\rho}}"

where B= bulk modulus , and "\rho" = density of the fluid

v= "\sqrt { \frac{2.8\times 10^{10}}{1.36 \times 10^4 }} =" 1434.8601 m/s